What is the moment of inertia for a 0.15-Kg billiard ball that has a radius of 0.1m?
1
Expert's answer
2012-03-22T12:11:21-0400
Solid ball of mass m and radius R, rotating around an axis which passes through the center, has the moment of inertia

I = 2/5·m·R².

So, the moment of inertia for a 0.15-Kg billiard ball that has a radius of 0.1m is

I = 2/5·0.15·0.1² = 3/5000 kg·m².